OpenStudy (anonymous):

if triangle abc is similar to triangle XYZ and side AB = 16 when side XY = 18, what is the length of side BC when side YZ = 27

OpenStudy (campbell_st):

well without a diagram it looks a bit like this when comparing ratios or corresponding sides \[\frac{AB}{XY} = \frac{BC}{YZ}\] so substitute the know values \[\frac{16}{18} = \frac{BC}{27}\] Hope this helps you find the answer